OpenStudy (anonymous):

Find the anti-derivative of F(x)=1/sqrt(x)

OpenStudy (anonymous):

\[ F(x) = x^{-1/2} \]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Now, remember that \[ (x^n) '= nx^{n-1} \]Which means: \[ \left(\frac{x^n}{n}\right)' = x^{n-1} \]

OpenStudy (anonymous):

Well... \[ \frac{1}{\sqrt x} = \frac{1}{x^{1/2}} = x^{-1/2} \]

OpenStudy (anonymous):

my textbook says the answer is 2t^1/2+c

OpenStudy (anonymous):

Now, remember that \[ (x^n) '= nx^{n-1} \]Which means: \[ \left(\frac{x^n}{n}\right)' = x^{n-1} \]If we let \(m=n-1\) then: \[ \left(\frac{x^{m+1}}{m+1}\right)' = x^{m} \]

OpenStudy (anonymous):

\[\int\limits_{ }^{ } x^n~dx~=\frac{n^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

I never found the anti derivative yet.

OpenStudy (anonymous):

2x*

OpenStudy (anonymous):

to what power?

OpenStudy (anonymous):

ok sorry wio...contin ue

OpenStudy (anonymous):

Anyway, since we want anti derivative of \[ x^{-1/2} \]We can use:\[ \left(\frac{x^{m+1}}{m+1}\right)' = x^{m} \]with \(m=-1/2\).

OpenStudy (anonymous):

This gives: \[ \left(\frac{x^{1/2}}{1/2}\right)' = x^{-1/2} \]

OpenStudy (anonymous):

So anti derivative is \[ 2\sqrt x+C \]

OpenStudy (anonymous):

\[\int\limits_{ }^{ } x^{-1/2}~dx=~\frac{x^{-1/2~~~+1}}{-1/2~~~~+1}=\frac{x^{1/2}}{1/2}=2x^{1/2}=2\sqrt{x}\color{red}{+C}\]

OpenStudy (anonymous):

thank you guys very much!